# Unifying Multiple Foundation Models for Advanced Computational Pathology

### v3

@@ 3 Shazam improves tile-level performance across diverse histopathology tasks
Tile-level tasks provide a stringent assessment of a model’s capacity to capture fine-grained morphological patterns, local microenvironmental variation, and institution- or stain-specific distribution shifts, all of which are typically underrepresented in whole-slide aggregation benchmarks. Rather than relying on a single foundation model and its fixed inductive biases, Shazam integrates multi-scale features from several complementary pathology foundation models and adaptively selects informative channels at inference time. This produces tile-level descriptors that reflect a coordinated ensemble of morphological primitives with greater flexibility across heterogeneous tissue contexts. To systematically evaluate this integrative design, we benchmarked Shazam across eleven diverse tile-level datasets: HunCRC for colorectal tissue classification[pataki2022huncrc], UniToPatho for colorectal polyp grading[barbano2021unitopatho], PCAM[bejnordi2017diagnostic] for metastatic detection, ESCA for esophageal carcinoma subtyping[tolkach_yuri_2023_7548828], PanCancer-TIL for tumor-infiltrating lymphocytes assessment[saltz2018spatial], CCRCC-TCGA-HEL[brummer2023computational] for renal carcinoma tissue classification, BACH for breast cancer categorization[aresta2019bach], CRC-MSI for microsatellite instability screening[kather2019deep], PanCancer-TCGA for pan-cancer tissue classification[komura2022universal], and Chaoyang for colon pathology[zhu2021hard]. These datasets span a wide range of tasks and clinical settings, providing a comprehensive evaluation of Shazam’s tile-level generalizability.
@@
Across the eleven tile-level benchmarks, Shazam achieved the strongest overall performance, with a mean ranking score of 1.36 compared with 2.46 for the strongest baseline, Virchow2, underscoring the consistent advantage of its integrative multi-model formulation (Supplementary Tables 7-17). Shazam delivered the highest F1 score on several representative datasets. On PCAM, it achieved 0.914 and surpassed Virchow2 by approximately 1.4%. On PanCancer-TIL, Shazam reached 0.852, outperforming Virchow2 by about 1.0%. On PanCancer-TCGA, it attained 0.844, exceeding H-optimus-1 at 0.833 by roughly 1.1%. These improvements illustrate how Shazam leverages complementary multi-scale features from multiple foundation models to produce more discriminative and robust tile-level descriptors across varied tissue contexts. Although Shazam led performance on most benchmarks, its advantage narrowed in datasets where a single baseline model exhibited a pronounced task-specific strength. For instance, on UniToPatho, Virchow2 achieved a F1 score of 0.820, noticeably higher than the other individual models, and Shazam obtained 0.803, reflecting that the ensemble can inherit attenuated signals when several constituent models underperform. Even in such settings, Shazam remained consistently competitive across metrics, maintaining reliable performance despite variations in stain, scanner, and tissue preparation. This stability highlights Shazam’s capacity to generalize fine-grained morphological patterns at the tile level, where subtle artifacts, microenvironmental cues, and local histologic variability pose the greatest challenges for computational pathology models.
@@ 4 Shazam improves pathology VQA
Visual Question Answering (VQA) aims to answer natural-language queries based on visual evidence and has emerged as a promising framework for assisting clinicians in retrieving task-specific diagnostic information from histopathology images. To examine Shazam’s capability in this setting, we used the PathVQA[he2020pathvqa] dataset, which is among the largest resources for pathology VQA. The dataset contains 32,799 image–question–answer triplets, divided into a training set of 16,400 samples, a validation set of 9,840 samples, and a test set of 6,560 samples. Overall performance across the evaluated pathology foundation models is presented in Fig. 4b, with detailed results provided in Supplementary Table 18.
Figure 4: a, Quantitative comparison of tile-level performance across eleven classification benchmarks, measured by the F1 score for each foundation model. The first panel summarizes the average ranking of all models across tasks, where Shazam achieved the lowest (best) mean rank. The p-value shown in this panel corresponds to a one-sided Wilcoxon signed-rank test evaluating whether Shazam’s rankings are significantly better than those of the strongest individual baseline. The remaining panels display the per-dataset distribution of F1 scores across repeated trials. b, Overall VQA accuracy for different pathology foundation models on the PathVQA test set. Error bars denote 95% confidence intervals obtained via 1000 times bootstrapping. Shazam achieves the highest accuracy among all evaluated models. c, A question and the answers generated by foundation models related to the query image.
@@
We further provide qualitative comparisons in Fig. 4c, where we illustrate a representative query image, question, and predicted answers across different foundation models. Shazam and UNI 2 consistently produced more reliable and clinically meaningful responses than the other evaluated models, supporting the quantitative advantages observed in the benchmark analysis.
@@ Discussion
Despite its strong empirical performance, Shazam has several limitations that warrant further investigation. Because Shazam requires all constituent foundation models to perform feature extraction during both training and inference, its computational cost grows approximately linearly with the number of integrated models. This requirement may limit its practicality in settings where low latency or constrained hardware resources are essential. Recent progress in developing lightweight pathology foundation models with substantially fewer parameters yet competitive performance provides a promising avenue for reducing this overhead[filiot2025distilling], and future versions of Shazam could incorporate such models to achieve a more favorable efficiency–performance balance. Another limitation lies in the breadth of external validation. Although Shazam demonstrates consistent gains across a wide range of internal benchmarks, the evaluation on independent datasets remains limited. Expanding external validation to additional cohorts and institutions will be important for fully establishing its robustness under real-world distributional variability. Continued exploration along these directions may further enhance the scalability, generalizability, and clinical readiness of Shazam.
@@ 1 Shazam
Shazam is a multi-scale fusion framework designed to synthesize visual representations from diverse pathology foundation models. It consists of four key components: (1) multi-scale teacher feature extraction, (2) dimensional alignment, (3) adaptive mixture-of-experts (MoE) fusion, and (4) task-specific online distillation. Rather than relying on a single pretrained model, Shazam leverages complementary information distributed across different architectures and across multiple semantic scales.
@@
For each foundation model, we extract visual features at three characteristic depths—early, middle, and final layers—corresponding to low-, mid-, and high-level semantic representations. To ensure reproducibility across diverse architectures, we determine the extraction points mathematically. Let $L$ denote the total number of transformer blocks in a given teacher model. We place non-invasive forward hooks at block indices defined by $l_{\text{low}}=\lfloor 0.33\times L\rfloor$, $l_{\text{mid}}=\lfloor 0.66\times L\rfloor$, and $l_{\text{high}}=L$. During the training of Shazam, all teacher models remain frozen to serve as stable feature extractors. This strategy offers a consistent and lightweight mechanism for capturing multi-scale information across the five foundation models used in this study, namely UNI 2[chen2024towards], Virchow2[zimmermann2024virchow2], H-optimus-1[hoptimus1], Prov-Gigapath[xu2024whole], and Phikon-v2[filiot2024phikonv2largepublicfeature].
@@
Because the teacher models operate in different embedding spaces, all extracted features are first mapped into a shared latent dimension using linear projection layers. At each semantic scale, Shazam then applies a Mixture-of-Experts (MoE)[fedus2022switch] module that learns to weight teacher-specific contributions. The MoE gating mechanism conditions on the concatenated teacher representations and produces an adaptive fusion, enabling the model to emphasize different visual experts depending on the input characteristics. The fused representations are refined through a small stack of self-attention blocks, which encourage information exchange across models. Higher-level fused features incorporate residual connections from preceding scales, allowing coarse structural cues from shallow layers to influence deeper semantic representations.
@@
For whole-slide image (WSI) analysis, each teacher produces patch-level features at all three scales. These patch embeddings are aggregated into WSI-level representations using the Attention-Based Multiple Instance Learning (ABMIL) mechanism proposed by Ilse et al.[ilse2018attention]. Shazam employs scale-specific, independent ABMIL heads to aggregate features at each of the low, mid, and high scales, preserving hierarchical distinctiveness before the final fusion and prediction. Tile-level tasks directly use per-image features without patch aggregation.
@@
To encourage the fused representation to preserve the strengths of individual foundation models, Shazam introduces a multi-level distillation objective. We apply $L_{2}$-normalization to both the student and teacher features before computing the distillation loss, which is essential for stable cosine distance calculation. This multi-scale approach encourages the student to maintain consistency with the teacher’s representations across different semantic hierarchies. For each scale $s$, the student feature $z_{s}$ is aligned with each teacher’s feature vector $t^{(i)}_{s}$ using a combination of normalized cosine distance and Huber loss:
$\mathcal{L}_{\text{distill}}^{(s,i)}=\left(1-\frac{z_{s}\cdot t^{(i)}_{s}}{\|z_{s}\|_{2}\|t^{(i)}_{s}\|_{2}}\right)+\mathrm{Huber}(z_{s},t^{(i)}_{s})$,
The overall distillation loss averages across all scales and all $N=5$ teachers:
$\mathcal{L}_{\text{distill}}=\frac{1}{3N}\sum_{s\in\{\mathrm{low},\mathrm{mid},\mathrm{high}\}}\sum_{i=1}^{N}\mathcal{L}_{\text{distill}}^{(s,i)}.$
@@ 3 Spatial Transcriptomic Experiments
Spatial transcriptomics measures gene expression while preserving the anatomical location of each measurement within intact tissue sections [staahl2016visualization]. By linking RNA abundance directly to histological context, these technologies enable spatially resolved characterization of cell states, signaling programs, and tumor–microenvironment interactions that cannot be inferred from morphology or bulk profiling alone [staahl2016visualization, pineiro2022research]. This integration of molecular and histologic information provides a powerful framework for understanding tissue organization and disease progression.
To define clinically meaningful prediction targets, we curated cohort-specific gene lists guided by FDA-recognized biomarkers in OncoKB [oncokb_fda_recognition]. Cohort-level preprocessing was then applied to reduce sparsity and ensure adequate expression signal. Slides were excluded if at least half of the selected genes had zero expression, and genes were excluded if they were zero in at least half of the slides within the same cohort. The resulting filtered datasets were used for all downstream analyses. Cohort-specific gene lists are summarized in Supplementary Table 1.
We conducted experiments using the publicly available HEST 1k dataset [jaume2024hest], which provides paired spatial transcriptomic profiles and matched H&E-stained whole-slide images across diverse organs. From this dataset, we selected eight organ-specific cohorts with sufficient paired samples: breast, prostate, skin, bowel, liver, kidney, brain, and bladder. To prevent patient-level data leakage, we followed a patient-stratified protocol: samples without patient identifiers were assigned exclusively to the training pool, and samples with identifiers were split at the patient level so that all spots and patches from each patient occurred within a single fold [jaume2024hest].
@@
4 Patch-level Classification

### v4

@@ 3 Shazam improves tile-level performance across diverse histopathology tasks
Tile-level tasks provide a stringent assessment of a model’s ability to capture fine-grained morphological patterns, local microenvironmental variation, and institution- or stain-specific distribution shifts, factors that are often attenuated in whole-slide aggregation benchmarks. Rather than relying on a single foundation model with fixed inductive biases, Shazam integrates multi-scale features from multiple complementary pathology foundation models and adaptively selects informative channels at inference time. This design yields flexible tile-level representations that combine diverse morphological primitives and better accommodate heterogeneity across tissues and acquisition settings.
To systematically evaluate this integrative capability, we benchmarked Shazam on eleven diverse tile-level datasets, including ESCA for esophageal carcinoma subtyping[42], PanCancer-TCGA for pan-cancer tissue classification[25], PanCancer-TIL for tumor-infiltrating lymphocyte assessment[37], Chaoyang for colon pathology[49], BACH for breast cancer categorization[3], CCRCC-TCGA-HEL for renal carcinoma tissue classification[10], CRC-MSI for microsatellite instability screening[24], PCAM for metastatic detection[8], UniToPatho for colorectal polyp grading[6], CRC-100K[23] and HunCRC for colorectal tissue classification[35].
As summarized in Fig. 4a, these benchmarks collectively span a broad spectrum of tile-level task categories, including normal tissue identification, precancerous lesion grading, malignant tumor recognition, tumor microenvironment characterization, and molecular phenotype prediction. This heterogeneous task coverage enables a comprehensive and clinically representative evaluation of Shazam’s tile-level generalization performance.
@@
Across the eleven tile-level benchmarks, Shazam achieved the strongest overall performance, with a mean ranking score of 1.36 compared with 2.46 for the strongest baseline, Virchow2, underscoring the consistent advantage of its integrative multi-model formulation (Fig. 4a; Supplementary Tables 717). Shazam delivered the highest F1 score on several representative datasets. On PCAM, it achieved 0.914 and surpassed Virchow2 by approximately 1.4%. On PanCancer-TIL, Shazam reached 0.852, outperforming Virchow2 by about 1.0%. On PanCancer-TCGA, it attained 0.844, exceeding H-optimus-1 at 0.833 by roughly 1.1%. These improvements illustrate how Shazam leverages complementary multi-scale features from multiple foundation models to produce more discriminative and robust tile-level descriptors across varied tissue contexts.
Figure 4: Evaluation of tile-level classification and pathology VQA. a, Coverage of tile-level task categories across benchmark datasets, with filled blocks indicating the availability of corresponding annotations. b, Quantitative comparison of tile-level performance across eleven classification benchmarks measured by F1 score. The reported $p$-value is from a one-sided Wilcoxon signed-rank test comparing Shazam with the strongest individual baseline; per-dataset F1 score distributions are shown across repeated trials. c, Overall VQA accuracy on the PathVQA test set, with error bars indicating 95% confidence intervals estimated by 1000 bootstrap samples. d, Example question and model-predicted answers for a representative query image.
Although Shazam led performance on most benchmarks, its advantage narrowed in datasets where a single baseline model exhibited a pronounced task-specific strength. For instance, on UniToPatho, Virchow2 achieved a F1 score of 0.820, noticeably higher than the other individual models, and Shazam obtained 0.803, reflecting that the ensemble can inherit attenuated signals when several constituent models underperform. Even in such settings, Shazam remained consistently competitive across metrics, maintaining reliable performance despite variations in stain, scanner, and tissue preparation. This stability highlights Shazam’s capacity to generalize fine-grained morphological patterns at the tile level, where subtle artifacts, microenvironmental cues, and local histologic variability pose the greatest challenges for computational pathology models.
@@ 4 Shazam improves pathology VQA
Visual Question Answering (VQA) aims to answer natural-language queries based on visual evidence and has emerged as a promising framework for assisting clinicians in retrieving task-specific diagnostic information from histopathology images. To examine Shazam’s capability in this setting, we used the PathVQA[19] dataset, which is among the largest resources for pathology VQA. The dataset contains 32,799 image–question–answer triplets, divided into a training set of 16,400 samples, a validation set of 9,840 samples, and a test set of 6,560 samples. Overall performance across the evaluated pathology foundation models is presented in Fig. 4c, with detailed results provided in Supplementary Table 19.
@@
We further provide qualitative comparisons in Fig. 4d, where we illustrate a representative query image, question, and predicted answers across different foundation models. Shazam and UNI 2 consistently produced more reliable and clinically meaningful responses than the other evaluated models, supporting the quantitative advantages observed in the benchmark analysis.
Figure 5: Ablation analysis of Shazam. a, Effect of teacher supervision on HEST-Bowel. Box plots show the median (center line), mean (rhombus),interquartile range (IQR, box limits represent 25th and 75th percentiles), and whiskers extending to 1.5 × IQR; outliers are shown as individual points. b, Complementarity of semantic feature representations for WSI-level survival prediction on TCGA-COADREAD. c, Contribution of the Mixture-of-Experts (MoE) module on CRC-100K and CCRCC classification tasks.
5 Ablation study of Shazam
To further dissect the sources of Shazam’s performance gains, we conducted a unified ablation study examining the contributions of teacher supervision, semantic feature representations, and adaptive aggregation. These experiments are designed to quantify how performance degrades as informative signals are removed and to clarify which components contribute most to robustness across tasks.
We first analyzed the role of teacher supervision by progressively removing teacher models on the HEST-Bowel dataset (Fig. 5a; Supplementary Table 20). Teachers were ablated in descending order of their standalone performance. With all teachers retained, Shazam achieved a mean PCC of 0.242. Removing the strongest teacher (UNI-v2) reduced the mean PCC to 0.213, corresponding to a relative decrease of 12.0%. In contrast, removing the second-best teacher (Virchow2) led to a smaller drop, with a mean PCC of 0.233 (a 3.7% decrease). Together, these results support our hypothesis that prioritizing task-aligned foundation models yields more effective supervision than treating all teachers equally, and that integrating such models within Shazam leads to measurable and robust performance improvements.
We next quantified the contribution of feature representations at different levels of semantic abstraction for whole-slide survival prediction on TCGA-COADREAD (Fig. 5b; Supplementary Table 21). Using features from a single level yielded C-index values in the range of 0.680–0.689. Pairwise combinations improved performance by up to 0.022 in absolute C-index, with the low- and high-level combination achieving 0.702. Aggregating low-, mid-, and high-level features achieved the highest C-index of 0.710, representing an absolute improvement of 0.030 over the best single-level configuration and 0.008 over the best pairwise combination. These results demonstrate that representations encoding local morphology, intermediate tissue organization, and global contextual patterns capture complementary prognostic information.
Finally, we evaluated the effect of the Mixture-of-Experts (MoE) module on CRC-100K and CCRCC-TCGA-HEL classification tasks (Fig. 5c; Supplementary Tables 22-23). On CRC-100K, incorporating MoE improved balanced accuracy from 0.955 to 0.960 (+0.005), weighted F1 from 0.951 to 0.956 (+0.005), and top-1 accuracy from 0.970 to 0.972 (+0.002). On CCRCC-TCGA-HEL, the corresponding improvements were smaller but consistent, with weighted F1 increasing from 0.973 to 0.975 (+0.002) and top-1 accuracy from 0.973 to 0.975 (+0.002). These results suggest that adaptive expert routing provides measurable but dataset-dependent gains over uniform feature aggregation.
Together, these ablation studies indicate that Shazam’s performance gains arise from the interaction of strong but uneven teacher supervision, complementary semantic feature representations, and adaptive aggregation. While the largest gains stem from retaining informative teachers and multi-level features, MoE further refines performance by enabling task-adaptive fusion of heterogeneous representations.
@@ Discussion
Despite its strong empirical performance, Shazam has several limitations that warrant further investigation. Because Shazam requires all constituent foundation models to perform feature extraction during both training and inference, its computational cost grows approximately linearly with the number of integrated models. This requirement may limit its practicality in settings where low latency or constrained hardware resources are essential. Recent progress in developing lightweight pathology foundation models with substantially fewer parameters yet competitive performance provides a promising avenue for reducing this overhead[17], and future versions of Shazam could incorporate such models to achieve a more favorable efficiency–performance balance. Another limitation lies in the breadth of external validation. Although Shazam demonstrates consistent gains across a wide range of internal benchmarks, the evaluation on independent datasets remains limited. Expanding external validation to additional cohorts and institutions will be important for fully establishing its robustness under real-world distributional variability. Continued exploration along these directions may further enhance the scalability, generalizability, and clinical readiness of Shazam.
@@ 1 Shazam
Shazam is a multi-scale fusion framework designed to synthesize visual representations from diverse pathology foundation models. It consists of four key components: (1) multi-teacher and multi-scale representation extraction, (2) Feature alignment and scale-wise Fusion, (3) Patch-to-slide aggregation for WSI tasks, and (4) Distillation across scales and teachers. Rather than relying on a single pathology foundation model, Shazam leverages complementary information distributed across different architectures and across multiple semantic scales.
@@
For each foundation model, we extract visual features at three characteristic depths—early, middle, and final layers—corresponding to low-, mid-, and high-level semantic representations. To ensure reproducibility across diverse architectures, we determine the extraction points mathematically. Let $L$ denote the total number of transformer blocks in a given teacher model. We place non-invasive forward hooks at block indices defined by $l_{\text{low}}=\lfloor 0.33\times L\rfloor$, $l_{\text{mid}}=\lfloor 0.66\times L\rfloor$, and $l_{\text{high}}=L$. During the training of Shazam, all teacher models remain frozen to serve as stable feature extractors. This strategy offers a consistent and lightweight mechanism for capturing multi-scale information across the five foundation models used in this study, namely UNI 2[12], Virchow2[50], H-optimus-1[9], Prov-Gigapath[48], and Phikon-v2[18].
@@
In Shazam, the multi-scale features from multiple teacher models are first aligned into a shared embedding space to enable effective fusion. Given the multi-scale features from different foundation models, we concatenate the features across all experts and apply a gating mechanism to weight the contribution of each teacher’s representation. Specifically, let $\mathbf{f}_{1},\mathbf{f}_{2},\dots,\mathbf{f}_{N}$ denote the feature representations from $N$ teacher models, where each $\mathbf{f}_{i}\in\mathbb{R}^{d_{\text{model}}}$ is the feature from the $i$-th teacher. The features from each teacher model are concatenated into a single vector:
$\mathbf{f}_{\text{concat}}=\text{concat}(\mathbf{f}_{1},\mathbf{f}_{2},\dots,\mathbf{f}_{N})\in\mathbb{R}^{N\cdot d_{\text{model}}}$
A gating network is applied to the concatenated feature vector to compute the contribution weights for each expert. The gating mechanism consists of a simple multilayer perceptron (MLP) that takes the concatenated features and produces the gating scores, $\mathbf{g}$. Specifically, the gating scores are computed as:
$\mathbf{g}=\text{softmax}\left(\text{MLP}(\mathbf{f}_{\text{concat}})\right)$
where $\mathbf{f}_{\text{concat}}$ is the concatenated feature vector from all teachers, and MLP represents a multi-layer perceptron. The output, $\mathbf{g}$, is a set of weights that determine the contribution of each expert’s feature to the final fused representation.
The weighted feature from each teacher is first computed as $g_{i}\mathbf{f}_{i}$, where $g_{i}$ denotes the gating score for the $i$-th teacher and $\mathbf{f}_{i}\in\mathbb{R}^{d}$ is the corresponding feature representation. Rather than concatenating these weighted features into a single vector, the Mixture-of-Experts (MoE) module organizes them into a feature matrix, with each row encoding the gated contribution of one teacher. Formally, the fused representation is defined as
$\mathbf{F}_{\text{fused}}=\begin{bmatrix}g_{1}\mathbf{f}_{1}^{\top}\\ g_{2}\mathbf{f}_{2}^{\top}\\ \vdots\\ g_{N}\mathbf{f}_{N}^{\top}\end{bmatrix}\in\mathbb{R}^{N\times d},$
where $N$ is the number of teacher models and each row corresponds to a teacher-specific, gated feature vector. This fused feature matrix preserves the individual contributions of different teachers and is subsequently passed to the self-attention layers, which model interactions across teachers and adaptively aggregate their information for downstream prediction.
To further enhance the quality of the fused representation, we apply a small stack of self-attention blocks to the stacked weighted features. These blocks facilitate information exchange across features at different semantic scales, improving the model’s ability to capture both fine-grained details and high-level context. The self-attention mechanism uses the stacked weighted features as input and learns to focus on important interactions between the features:
$\mathbf{f}_{\text{att}}=\text{SelfAttention}(\mathbf{f}_{\text{fused}})$
Finally, the output from the self-attention block is passed through a layer normalization [4] step to stabilize the final feature representation:
$\mathbf{f}_{\text{final}}=\text{LayerNorm}(\mathbf{f}_{\text{att}})$
This multi-scale, expert-weighted fusion process enables Shazam to capture complementary information from multiple teacher models, facilitating effective feature alignment and fusion across scales.
@@
For whole-slide image (WSI) analysis, each teacher produces patch-level features at all three scales. These patch embeddings are aggregated into WSI-level representations using the Attention-Based Multiple Instance Learning (ABMIL) mechanism proposed by Ilse et al.[20]. Shazam employs scale-specific, independent ABMIL heads to aggregate features at each of the low, mid, and high scales, preserving hierarchical distinctiveness before the final fusion and prediction. Tile-level tasks directly use per-image features without patch aggregation.
@@
To ensure that the fused representation retains the complementary strengths of individual foundation models, Shazam employs a multi-level distillation objective across three semantic scales. For each scale $s\in\{\mathrm{low},\mathrm{mid},\mathrm{high}\}$, we $\ell_{2}$-normalize both the student feature $z_{s}$ and the $i$-th teacher feature $t_{s}^{(i)}$ prior to distillation to stabilize cosine-based alignment. The student is then supervised to match each teacher using a combination of cosine distance and an element-wise Huber loss:
$\mathcal{L}_{\mathrm{distill}}^{(s,i)}=\left(1-\frac{\langle z_{s},\,t_{s}^{(i)}\rangle}{\|z_{s}\|\,\|t_{s}^{(i)}\|}\right)+\mathrm{Huber}\!\left(z_{s},\,t_{s}^{(i)}\right),$ (1)
where $\langle\cdot,\cdot\rangle$ denotes the inner product and $\|\cdot\|$ denotes the Euclidean norm.
Element-wise Huber loss. Let $e_{s,k}^{(i)}=z_{s,k}-t_{s,k}^{(i)}$ be the discrepancy at feature dimension $k\in\{1,\ldots,d\}$. We define the element-wise Huber loss as
$\mathrm{Huber}\!\left(z_{s},t_{s}^{(i)}\right)=\frac{1}{d}\sum_{k=1}^{d}\begin{cases}\frac{1}{2}\left(e_{s,k}^{(i)}\right)^{2},&\left|e_{s,k}^{(i)}\right|\leq\delta,\\[4.0pt] \delta\left(\left|e_{s,k}^{(i)}\right|-\frac{1}{2}\delta\right),&\left|e_{s,k}^{(i)}\right|>\delta,\end{cases}$ (2)
where $\delta>0$ is the transition threshold and $d$ is the feature dimension. This formulation is quadratic for small discrepancies and linear for larger deviations, reducing sensitivity to outlier channels in teacher features.
Finally, we average the distillation objective over all scales and all $N=5$ teacher models:
$\mathcal{L}_{\mathrm{distill}}=\frac{1}{3N}\sum_{s\in\{\mathrm{low},\mathrm{mid},\mathrm{high}\}}\sum_{i=1}^{N}\mathcal{L}_{\mathrm{distill}}^{(s,i)}.$ (3)
@@ 3 Spatial Transcriptomic Experiments
Spatial transcriptomics measures gene expression while preserving the anatomical location of each measurement within intact tissue sections [40]. By linking RNA abundance directly to histological context, these technologies enable spatially resolved characterization of cell states, signaling programs, and tumor–microenvironment interactions that cannot be inferred from morphology or bulk profiling alone [40, 36]. This integration of molecular and histologic information provides a powerful framework for understanding tissue organization and disease progression.
We conducted experiments on the publicly available HEST-1k dataset [22], which collected paired spatial transcriptomic profiles and matched H&E-stained whole-slide images across multiple organs. Cohorts were derived from the dataset metadata (CSV) by restricting to malignant cases (disease_state = cancer) and selecting organ labels in breast, prostate, skin, bowel, liver, kidney, brain, bladder, yielding eight organ-specific subsets with sufficient paired samples. To avoid patient-level leakage, we used a patient-stratified split: samples lacking patient identifiers were assigned to the training set only, whereas samples with identifiers were partitioned at the patient level such that all associated spots and image patches from a given patient appeared in a single fold.
To define clinically meaningful prediction targets, we curated cohort-specific gene lists guided by FDA-recognized biomarkers in OncoKB [33]. Cohort-level preprocessing was then applied to reduce sparsity and ensure adequate expression signal. Slides were excluded if at least half of the selected genes had zero expression, and genes were excluded if they were zero in at least half of the slides within the same cohort. The resulting filtered datasets were used for all downstream analyses. Cohort-specific gene lists are summarized in Supplementary Table 1 and Supplementary Table 2.
@@
4 Survival Analysis
In oncology, prognostic modeling aims to link patient time-to-event outcomes (e.g., progression and overall survival) to measurable evidence. By providing a wealth of tissue-scale data, whole-slide images allow survival modeling to take advantage of histomorphological cues and connect them to clinical prognosis.
[…]
In this study, we adopt the Attention-Based Multiple Instance Learning (ABMIL) framework [20] for survival analysis, employing the Negative Log-Likelihood (NLL) loss. we introduce a distillation loss term for model compression. Specifically, we employ Shazam student model, distilled from a larger teacher model, using a weighted distillation loss with $\lambda=0.01$ to align the student’s predictions with the teacher’s. Optimization is performed using AdamW with a learning rate of $2\times 10^{-4}$ and weight decay of $1\times 10^{-3}$.The dataset partitioning strategy follows that described in the GPFM study [28]. To assess the impact of different pathology foundation models on survival prediction, we perform 5-fold cross-validation and use the C-index as the evaluation metric.
Experiments are conducted on 10 TCGA datasets: TCGA-BRCA, TCGA-BLCA, TCGA-CESC, TCGA-COADREAD, TCGA-GBMLGG, TCGA-KIRC, TCGA-LUAD, TCGA-LUSC, TCGA-SKCM and TCGA-STAD. The results are reported in Supplementary Table 5- 7.
Breast Cancer Survival Prediction.
In the breast cancer survival prediction task, the TCGA-BRCA dataset [46] is adopted, containing 1,023 cases with a total of 1,089 WSIs. Following the GPFM partitioning strategy, we stratify survival times into four equally populated bins for both training and testing to ensure comparable censoring patterns across splits. For each fold of the 5-fold cross-validation, we perform an 8:2 label-stratified split, resulting in 821 training cases and 202 testing cases. The results across the 5 folds are presented in Supplementary Table 5.
Bladder Cancer Survival Prediction.
For bladder cancer survival prediction, we use the TCGA-BLCA dataset[46], containing 376 cases (446 WSIs). Using the same stratified binning and partitioning strategy as BRCA, each fold allocates 305 cases to training and 71 cases to testing. The results across the 5 folds are presented in Supplementary Table 5.
The TCGA-CESC[46] dataset includes 250 cases (260 WSIs). Using the same strategy, we assign 203 cases to training and 47 cases to testing per fold. The results across the 5 folds are presented in Supplementary Table 5.
Colon & Rectum Adenocarcinoma Survival Prediction.
We combine TCGA-COAD[46] (426 cases, 434 WSIs) and TCGA-READ (153 cases, 154 WSIs) into a single dataset of 579 cases (588 WSIs). Each fold assigns 464 cases to training and 115 to testing. The results across the 5 folds are presented in Supplementary Table 6.
Glioma Survival Prediction.
We merge TCGA-GBM (370 cases, 823 WSIs) and TCGA-LGG[46] (460 cases, 778 WSIs) into a combined glioma dataset of 830 cases (1,601 WSIs). Per fold, 667 cases are used for training and 163 for testing. The results across the 5 folds are presented in Supplementary Table 6.
Kidney Renal Clear Cell Carcinoma Survival Prediction.
The TCGA-KIRC dataset[46] comprises 498 cases (504 WSIs). Following the same partitioning strategy, each fold assigns 401 cases to training and 97 cases to testing. The results across the 5 folds are presented in Supplementary Table 6.
Lung Adenocarcinoma Survival Prediction.
For TCGA-LUAD[46], we use 455 cases (518 WSIs), splitting each fold into 366 training cases and 89 testing cases. The results across the 5 folds are presented in Supplementary Table 7.
Lung Squamous Cell Carcinoma Survival Prediction.
The TCGA-LUSC dataset[46] comprises 452 cases (484 WSIs). Using the same stratified splitting approach, 365 cases are assigned to training and 87 to testing in each fold. The results across the 5 folds are presented in Supplementary Table 7.
Skin Cutaneous Melanoma Survival Prediction.
The TCGA-SKCM dataset[46] contains 415 cases (456 WSIs). Each fold uses 337 cases for training and 78 for testing. The results across the 5 folds are presented in Supplementary Table 7.
Stomach Adenocarcinoma Survival Prediction.
The TCGA-STAD dataset[46] contains 363 cases (389 WSIs). In each fold, 293 cases are allocated to training and 70 to testing using the same stratified splitting procedure. The results across the 5 folds are presented in Supplementary Table 6.
5 Tile-level Classification